OpenStudy (anonymous):

I've solved all of these problems already except for 4.. I'm really stuck but I would seriously appreciate it if the answers were provided so i could see if my work is right. I'm including the answers I got.

OpenStudy (anonymous):

\[3(9^{x})-3^{x}=0\] \[e ^{2x}+7e ^{x}=-12\] \[\log_{4}x+\log_{4}(2x+1)=2 \] \[\ln(5x+1)=\ln(2x)+\ln2\] \[\log_{3}(x ^{2} )+\log_{9} (3)=\log_{27}(x) \]

OpenStudy (anonymous):

my answers were: 1. -1 2. No Solution 3. 2.589 4. I got stuck on this one 5. 0.7192

OpenStudy (anonymous):

\[\Large e^{\ln(5x+1)} = e^{\ln(2x) + \ln(2)}\] \[\Large 5x+1 = e^{\ln(2x)}e^{\ln(2)}\] \[\Large 5x+1 = 2x*2\] Understand?

OpenStudy (anonymous):

So that way they cancel right?

OpenStudy (anonymous):

So x is -1?

OpenStudy (anonymous):

Good.

OpenStudy (anonymous):

Number 3 is correct, but you should have gotten multiple solutions.

OpenStudy (anonymous):

Yeah one was negative but you can't have a negative answer.

OpenStudy (anonymous):

Ah, good call =)

OpenStudy (anonymous):

Extraneous solution.

OpenStudy (anonymous):

all the rest of answers right?